# Machine-learning Guided Search for Phonon-mediated Superconductivity in Boron and Carbon Compounds

### v2

@@ II.6 Imaginary phonon modes and superconductivity
For the compounds with imaginary phonon modes away from the point, i.e. MoB2-type, we also stabilized these dynamically unstable compounds with larger electronic smearing of 0.1 Ry and tabulated the results in Table. 3. For example, compounds like MoB2 in the MgB2 structure have shown experimentally measured Tc of 32 K under high pressure around 110 GPa [34]. Another notable example in Table 3 is Ca5B3N6 with of 0.03 eV/atom, which exhibits dynamical instability at the point and displayed a significant mode-resolved near its unstable phonon modes, as depicted in Fig. 7 (d)-(e). After stabilizing the system with electronic smearing of 0.06 Ry, Ca5B3N6 exhibits a of 1.5. The is found to be 372 K. Moreover, when computing (with broadening parameter = 0.01 Ry), the Tc is determined to be 35 K with Coulomb potential =0.16 and 42.4 K with =0.10, much larger than that of MgB2 (16-20 K) computed from isotropic approximation. It should be noted that Ca5B3N6 has been synthesized in cubic structure and Im3′m (229) SG with partial occupancy in the 8c site of Ca [172]. The crystal structure of this boronitride (Fig. 7(a)-(c)) has a cage-like structure, similar to XB3C3 borocarbides (X Ca,Ba,Sr,Y,La). For the stoichiometric Ca5B3N6, Fig. 7(f) and (g) present the isotropic Eliashberg spectral function and electronic band structure, respectively. It shows an electron-doped band structure that connects to strong EPC as also found in other predicted SC compounds and studies. As listed in Table III, interestingly some of the trigonal compounds of NbMoC2 are in the same structure as the stable TaNbC2 in Table 1. This shows that in this particular structure, with the substitution of neighboring group of early TMs, although bringing dynamical instability, the phonons once stablized can still provide a large EPC contribution for a sizable Tc. As also listed in Table III, other compounds with sizable predicted Tc after stabilization that also near GS hull are some notable ternary Ru compounds, MoRuB2 at 15.6 K, RuVB2 at 15.0 K, Pd3CaB at 7.0 K and RuSc3C4 at 6.6 K.
@@ IV.1 Convergence with respect to Brillouin zone sampling (k-point mesh)
To identify cases of convergence failure (i.e., incorrect prediction of SC/NSC) related to Brillouin zone sampling, we analyzed the variation of Tc with respect to Gaussian broadening () in the double delta integration and developed a simple ansatz based on the converged results of MgB2, as described in the ”Convergence ansatz” section of the Supplemental Material (SM). This ansatz involves extracting Tc, similar to MgB2, and estimating the decay parameter (A) in the exponential variation of Tc with ,
@@ VI Acknowledgements
We thank Dr. Paul C. Canfield for the funding support, initiating the idea of searching for new phonon-mediated superconductors among boron and carbon compounds, and the helpful discussion throughout the project. This work was supported by Ames National Laboratory LDRD and U.S. Department of Energy, Office of Basic Energy Science, Division of Materials Sciences and Engineering. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358.
VII Author contributions
@@
VIII Competing interests
@@
X Statistics for compounds with unknown Tc
@@
XI Theory of superconductivity: Isotropic Approximation
@@
For computational feasibility, these Dirac deltas can be approximated by Gaussian functions with a broadening parameter [3], and EPC strength is given by Eqs.5 and 6 can be redefined as
@@
XII Convergence tests for MgB2 and AlB2
The ground-state total energy and phonon frequencies at -point are well converged for MgB2 with k-grids of 886, compared to the dense grid of 242424. K-mesh grid in materials project database for MgB2 is 887. Since the superconducting properties of MgB2 with k-grid of 888 doesn’t follow the trend of converged result of denser k-mesh with , we chose 886 so that we can use sufficient q-grids of 443 instead of using 999 k-grid and 333 q-grid. We utilized 886 k-grid and its multiple to compute the decay parameter. In this work, we are taking 0.75 obtained from solving anisotropic Migdal-Eliashberg equation using 0.16 as a reference [Comput. Phys. Commun. 209, 116 (2016)]. One can obtain this converged with denser k- and q- grids and , as shown in Figs. S2 S4. At last, we check whether the result is transferable to systems with larger unit cell where essentially smaller k-grids provide converged ground-state properties and q-grid, taken half of k-grid, sometime shrinks only to include the -point. To do that, we performed EPC calculations with 222 supercell of MgB2 which has 24 atoms per cell, and present the results in Fig. S5 and S6. Results confirm that q-grid as half as that of the k- grid can provide converged results even for larger systems. Instead of increasing coarse k- and q- grids, one can also utilize extremely large k- grid for interpolating EPC matrix element to achieve convergence, as shown in Figs. S4 and S7 respectively for MgB2 and AlB2. However, it also increases computational complexity for larger systems.
@@ XIII Convergence ansatz
Our investigation revealed that a reasonable number of calculations did not converge, leading to inaccurate predictions, while utilizing a coarse k-grid from MP database and fine grid for interpolation only twice that of the k-grid [Table ST3]. For example, AlB2 was predicted to be a superconductor (Tc 11 K) with the coarse grid [k-grid: 888, q-grid: 444, fine-k-mesh: 161616] from MP database, whereas a denser grid twice the size of the coarse grid corrected this inaccuracy and predicted AlB2 to be a nonsuperconductor, consistent with experimental observations [Fig. S7]. To identify such cases of convergence failure, we examined the variation of Tc with respect to and developed a simple ansatz based on the converged results of MgB2, as depicted in Fig. S9. The ansatz involves extracting Tc with a fixed broadening value (as in the case of MgB2) and estimating the decay parameter (A) in the exponential variation of Tc with as T. Unconverged results exhibit larger values of A, which decrease with a denser k-mesh.
@@
with also depends on . Despite of dependency of on EPC matrix () and phonon frequency () terms, the variation in is largely guided by the variation in or the Gaussian term G. Fig. S8(a)(a) shows the variation of G with respect to smearing for different values of , assuming independent of . The effect of on G dictates the variation of and hence Tc, with depends on the materials as well as on size of the grid to compute double delta integration. A larger represents more states contributing towards the double delta summation within space around the Fermi-level spanned by . In other word, will increase with if the density of states has local maximum close to the Fermi-level, and the variation of EPC properties will be similar to the Gaussian plots corresponding to 0.05. For Niobium, the variation of EPC properties with respect to follows similar to blue-dashed curve corresponding to = 0.01 [3], whereas for MgB2, the variation follows the Gaussian corresponding to 0.01. For and , if the Gaussian function recovers the delta function, while for the Gaussian drops to zero similar to 0.01 cases. Our primary focus is more on exponentially decaying cases.
@@
where is the Debye cutoff energy and U is electron-phonon coupling potential. One can established the relation between or on as and with constant. Here parameter is a constant with a positive value if decreases with increasing and becomes negative otherwise. Here A and B are constants to be determined with A denoting the coefficient of exponential increase (for negative A) or decrease (for positive A) of with respect to . To determine the value of these constants, we perform linear-fit of vs for different k- grids for MgB2, which is frequently studied both theoretically as well as experimentally and often challenging to obtain the converged SC properties [7, 8, 9]. Furthermore, one can capture the behavior of Tc vs for a wide range of Tc using MgB2, as shown in Fig. S8(a)(b). We utilize a smaller k-mesh of 664 and a slightly larger k′-grid of 886 for comparison [Fig. S8(a)(b) and (c)].
@@
Fig. S8(a) (b) represents the variation of with respect to , while Fig. S8(a) (c) presents vs plot. It exhibits nearly linear behavior up to , i.e. = 0.37, after which it deviates from linearity. The linear-fit has optimal coefficient of determination score (R2-score) for on vs data for k(charge density and EPC)-q(Phonon and EPC)-2k(interpolating EPC matrix) grids (Fig. S8(a) (c)) [Table within the Fig. S8(a)(c)]. Therefore, we chose in this work with R2-score of 0.9987. Besides different choice of leads to different values of A and B, it doesn’t have a significant role. For coarse grid, the Tc has a larger dependency on and shows larger exponential decay, compared to more converged calculations on denser grids. The exponential decay parameter “A” decreases from 19 to 10 with grids changing from coarse to dense one. This analysis suggests that A = 1213 can be used as a cutoff for this work, the calculations can be considered unconverged for A A, while the calculations can be considered converged for A A for accurate predictions. Fig. S8(a)(d) shows the convergence of the Tc with respect to q-point mesh, keeping k-grid fixed at 121212. A q-point grid as half as that of the ground-state k-point grid is sufficient to provide the converged results for MgB2. Fig. S8(a) (e) represents the spectral functions for various grids with k=886. Denser grid (2k) slightly blue shifts the spectral function peaks at lower frequency range (15.5-16.5 THz), while peaks at higher frequency ranges (20-22.5 THz) remain unaffected. This results slight change in from 0.78 to 0.72 and Tc changing from 20 K to 15-16 K [Figs. S1-S5]. A of 0.78 agrees with previous theoretical value of 0.75 from anisotropic Migdal-Eliashberg calculation using 0.16 [5], 0.71 from fully anisotropic SCDFT [10], and 0.73 from previous isotropic Eliashberg approximation [11] at slightly larger 0.015 Ry. This indicates a k-grid obtained from MP database with Ry already provide converged results in the case of MgB2. However, it is not always the case for other materials. Based on these results, parameters = 0.01 Ry with seems to be reasonable choice for smearing with 8 8 6 to 16 16 12 k-grids for MgB2, and for other systems for the sake of comparison. These parameters could also depend on pseudopotentials. In order to achieve converged results, it is necessary to use denser k- and q-grids. Subsequently, a double-delta integration should be performed, selecting the results that correspond to the limit of . Other details convergence tests of MgB2 with respect to Brillouin-zone sampling are presented in Figs. S2-S7.
@@
In Fig. S8(b), we show two cases of YB12 and C2I2Y2 in which the coarse k- grids from Materials Project and fine k- grid only twice of that result to qualitatively inaccurate results, unusually high Tc for the former while predicting negligible Tc for the latter. With coarse grids, the exponential decay parameter of YB12 is around 38, much larger than the critical value estimated for MgB2 (A 12). Increasing k-mesh grids two times in each direction, Tc drops from 40.6 K to 1.67 K at 0.01 Ry and drops to almost zero for 0.01 Ry indicating non-superconductor, which is in good agreement with the experiment [12]. We found a couple of cases such as C2I2Y2 when first decreases and increases later with (we set A = 0 in Fig. S10). Such behavior hasn’t been observed in G(,) vs plots for a wide range of (Fig. 8(a) of the main text). The coarse grids result to a very low Tc of 0.97 K for C2I2Y2. However, the denser mesh corrects the behavior with an exponential decay parameter of 12.5 much closer to the reference value of MgB2. Also, the critical temperature Tc improves to 5.29 K agreeing much more strongly to the experiment [13]. Note that, this analysis only works for compounds with non-zero for a wide range of . For 1 K at Ry (computational details for choosing Ry), 0.1 K for Ry, and rapidly drops to zero for larger , we identify the compound to be non-superconducting with the value of A as zero (A = 0 A, already converged). To summarize the ansatz, we present a simple schematics of the process shown as in Fig. S9.
@@
We present a few qualitatively inaccurate predictions with A A and improved results with a denser grid in Table ST3, while Table ST4 shows unaffected results (qualitatively) with denser grids for the cases with A A. For example, AlB2 is predicted to be superconductors with a coarse grid, while it changes to a non-superconductor with a denser grid, consistent with the experimental prediction. Similarly, of superconductors such as CaNiBN, C2I2Y2, and C2Cl2Y2 change from below 1 K (non-superconductor) to a larger value, and show a better agreement with the experiment. On the contrary, the qualitative results remain unchanged for converged calculations represented by A A, regardless of their agreement with the experimental references.
XIV Application of convergence ansatz on known compounds of SuperCon

### v3

@@ II.6 Imaginary phonon modes and superconductivity
For the compounds with imaginary phonon modes away from the point, i.e. MoB2-type, we also stabilized these dynamically unstable compounds with larger electronic smearing of 0.1 Ry and tabulated the results in Table. 3. For example, compounds like MoB2 in the MgB2 structure have shown experimentally measured Tc of 32 K under high pressure [34]. Moreover, we demonstrate that large electronic smearing can play a role similar to pressure in stabilizing imaginary phonon modes. The pressure-induced stabilization of MoB2 is presented in Supplementary Fig. S4. Our calculated structural phase transition from R-3m to P6/mmm phase at 75 GPa agrees well with the experimentally reported critical pressure of 70 GPa [34]. The calculated electron–phonon coupling (EPC) at 75 and 90 GPa yields Tc values of 28.4 K and 28.6 K, respectively, which are consistent with the experimental value of 32 K, as well as with results obtained from the electronic smearing approach used for high-throughput calculations. Notably, both approaches give comparable EPC parameters for the same stabilized low-lying phonon modes: at 75 GPa we obtain = 1.85, = 244 K, and Tc = 28.4 K, while the smearing-based approach (Table III) gives = 2.12, = 209 K, and Tc = 27.3 K. Another notable example in Table 3 is Ca5B3N6 with of 0.03 eV/atom, which exhibits dynamical instability at the point and displayed a significant mode-resolved near its unstable phonon modes, as depicted in Fig. 7 (d)-(e). After stabilizing the system with electronic smearing of 0.06 Ry, Ca5B3N6 exhibits a of 1.5. The is found to be 372 K. Moreover, when computing (with broadening parameter = 0.01 Ry), the Tc is determined to be 35 K with Coulomb potential =0.16 and 42.4 K with =0.10, much larger than that of MgB2 (16-20 K) computed from isotropic approximation. It should be noted that Ca5B3N6 has been synthesized in cubic structure and Im3′m (229) SG with partial occupancy in the 8c site of Ca [172]. The crystal structure of this boronitride (Fig. 7(a)-(c)) has a cage-like structure, similar to XB3C3 borocarbides (X Ca,Ba,Sr,Y,La). For the stoichiometric Ca5B3N6, Fig. 7(f) and (g) present the isotropic Eliashberg spectral function and electronic band structure, respectively. It shows an electron-doped band structure that connects to strong EPC as also found in other predicted SC compounds and studies. As listed in Table III, interestingly some of the trigonal compounds of NbMoC2 are in the same structure as the stable TaNbC2 in Table 1. This shows that in this particular structure, with the substitution of neighboring group of early TMs, although bringing dynamical instability, the phonons once stablized can still provide a large EPC contribution for a sizable Tc. As also listed in Table III, other compounds with sizable predicted Tc after stabilization that also near GS hull are some notable ternary Ru compounds, MoRuB2 at 15.6 K, RuVB2 at 15.0 K, Pd3CaB at 7.0 K and RuSc3C4 at 6.6 K.
@@ IV.1 Convergence with respect to Brillouin zone sampling (k-point mesh)
To identify cases of convergence failure (i.e., incorrect prediction of SC/NSC) related to Brillouin zone sampling, we analyzed the variation of Tc with respect to Gaussian broadening () in the double delta integration and developed a simple ansatz based on the converged results of MgB2, as described in the “Convergence tests for MgB2 and AlB2 (Supplementary Figures S5S10, Supplementary Tables S5S6)” and ”Convergence ansatz (Supplementary Figures S11S14, Supplementary Tables S7S9)” sections of the Supplemental Material (SM). This ansatz involves extracting Tc, similar to MgB2, and estimating the decay parameter (A) in the exponential variation of Tc with ,
@@ IV.2 Computational details
IV.3 Limitations
When evaluating the significance of machine-learning (ML) models trained on ab-initio data, it is essential to recognize both their capabilities and inherent limitations. First, the performance of any ML model is strongly dependent on the quality, diversity, and representativeness of its training dataset. Limited or biased datasets, which are common in computational materials science, can restrict a model’s predictive accuracy and transferability. Second, ML models are susceptible to overfitting, particularly when trained on small or noisy datasets, although careful validation procedures during model development can mitigate this issue to some extent. Third, the fidelity of ML predictions is ultimately bounded by the accuracy of the underlying ab-initio methods used to generate the training data. For example, the widely used isotropic Eliashberg approximation in superconductivity calculations neglects the inherently anisotropic and multi-band character of electron–phonon coupling. Although fully solving the anisotropic Migdal-Eliashberg equations (e.g., using EPW) yield more accurate estimates of superconducting properties, they are not computationally practical for large scale data generation. Consequently, ML models trained on data obtained from the isotropic Eliashberg approximation can function as efficient screening tools to identify promising candidate materials, which may then be examined with more accurate, anisotropic, and computationally demanding methods.
@@ VI Acknowledgements
We thank Dr. Paul C. Canfield for the funding support, initiating the idea of searching for new phonon-mediated superconductors among boron and carbon compounds, and the helpful discussion throughout the project.
VII Funding
This work was supported by Ames National Laboratory LDRD and U.S. Department of Energy, Office of Basic Energy Science, Division of Materials Sciences and Engineering. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358.
VIII Author information
VIII.1 Authors and Affiliations
Ames National Laboratory, Ames, Iowa 50011, USA
Niraj K. Nepal, Lin-Lin Wang
Department of Physics and Astronomy, Iowa State University, Ames, Iowa 50011, USA
Lin-Lin Wang
VIII.2 Contributions
@@
VIII.3 Corresponding author
Correspondence to Niraj K. Nepal (nepalneeraz@gmail.com)
IX Competing interests
@@
Niraj K. Nepal1 Lin-Lin Wang1,2
January 11, 2026
XI Statistics for compounds with unknown Tc
@@
| Training size | CGCNN | ALIGNN | ||||||
| (k) | Tc (K) | T (K) | (K) | Tc (K) | T (K) | |||
| 250 (Run1) | 0.15 | 72 | 1.8 | 1.3 | 0.15 | 56 | 1.7 | 2.2 |
| 323 (Run2) | 0.27 | 85 | 6.4 | 5.5 | 0.16 | 83 | 3.0 | 3.3 |
| 351 (Run3) | 0.31 | 70 | 3.9 | 6.5 | 0.24 | 60 | 3.0 | 4.4 |
| Model | EPC | Epoch | Best MAE | -score |
| CGCNN-250 | 48 | 0.145 | 0.75 | |
| CGCNN-250 | 60 | 49 | 0.45 | |
| CGCNN-250 | Tc | 13 | 2.80 | 0.56 |
| CGCNN-323 | 1190 | 0.203 | 0.66 | |
| CGCNN-323 | 590 | 57 | 0.5 | |
| CGCNN-323 | Tc | 736 | 2.73 | 0.31 |
| CGCNN-351 | 64 | 0.204 | 0.62 | |
| CGCNN-351 | 62 | 51 | 0.51 | |
| CGCNN-351 | Tc | 46 | 3.34 | 0.263 |
| ALIGNN-250 | 671 | 0.09 | 0.58 | |
| ALIGNN-250 | 672 | 27 | 0.47 | |
| ALIGNN-250 | Tc | 760 | 1.32 | 0.23 |
| ALIGNN-323 | 483 | 0.196 | 0.64 | |
| ALIGNN-323 | 293 | 57 | 0.63 | |
| ALIGNN-323 | Tc | 1275 | 2.33 | 0.32 |
| ALIGNN-351 | 186 | 0.22 | 0.72 | |
| ALIGNN-351 | 701 | 53 | 0.77 | |
| ALIGNN-351 | Tc | 617 | 2.85 | 0.38 |
| Training/test size | Property | Tc 1 | Tc 1 | ||
| CGCNN | ALIGNN | CGCNN | ALIGNN | ||
| 0.28 | 0.21 | 0.20 | 0.32 | ||
| 323/58 | 56 | 54 | 188 | 111 | |
| Tc | 3.60 | 1.98 | 2.23 | 4.30 | |
| 0.39 | 0.20 | 0.27 | 0.39 | ||
| 351/66 | 125 | 99 | 69 | 63 | |
| Tc | 4.67 | 1.97 | 6.23 | 5.95 | |
XII Classification models
In addition to regression, we also employ machine learning models for binary classification of materials into two categories: SC or NSC. We utilized the similar settings (batch, epochs, etc in config.json) as that of regression model training. Instead of finding a threshold Tc that maximizes the accuracy metrics, we simply perform the classification ML for a fix threshold Tc (0.5, 1.0, 2.0 K) and assess the model performance. The accuracy metrics used in our evaluation are presented in Table S4, and they are defined as follows:
Accuracy, Precision, Recall, and F1-score are performance metrics that evaluate different aspects of a classifier’s effectiveness. Accuracy measures the overall correctness of the classifier’s predictions. Precision quantifies the correctness of positive predictions, while Recall (sensitivity) assesses the successful identification of positive instances. The F1-score is the harmonic mean of Precision and Recall, providing a balance between the two. TP, TN, FP, and FN represent true positive (correctly identifying true superconductors), true negative (correctly identifying true nonsuperconductors), false positive (misidentifying nonsuperconductors as superconductors), and false negative (misidentifying superconductors as nonsuperconductors), respectively. The accuracy metrics of classification models for both Run 1 and Run 3 are displayed in Table. S4. Across all these metrics, the overall accuracies are reasonable, indicating a useful classification model despite the small dataset. By employing a classification threshold of Tc = 1 K, both ML models exhibit comparable performance. Notably, one can observe the enhanced accuracy of Run 3 in contrast to Run 1. This improvement might be attributed to the expanded training dataset size and the balanced representation of both SC and NSC instances after adding stabilized high Tc results. The ML model, when using a Tc threshold of 1 K, achieves an overall accuracy ranging from 60% to 70%, precision ranging from 70% to 90%, and an F1-score ranging between 50% and 60%. The ML models demonstrate lower recall values, ranging from 50% to 55%. Considering the data imbalance between the number of SC instances (approximately 40%) and NSC instances (approximately 60%), these recall values are reasonable.
| Run 1 | Run 3 | |||||||
XIII Theory of superconductivity: Isotropic Approximation
@@
For computational feasibility, these Dirac deltas can be approximated by Gaussian functions with a broadening parameter [3], and EPC strength is given by Eqs.9 and 10 can be redefined as
@@
XIV Convergence tests for MgB2 and AlB2
The ground-state total energy and phonon frequencies at -point are well converged for MgB2 with k-grids of 886, compared to the dense grid of 242424. K-mesh grid in materials project database for MgB2 is 887. Since the superconducting properties of MgB2 with k-grid of 888 doesn’t follow the trend of converged result of denser k-mesh with , we chose 886 so that we can use sufficient q-grids of 443 instead of using 999 k-grid and 333 q-grid. We utilized 886 k-grid and its multiple to compute the decay parameter. In this work, we are taking 0.75 obtained from solving anisotropic Migdal-Eliashberg equation using 0.16 as a reference [Comput. Phys. Commun. 209, 116 (2016)]. One can obtain this converged with denser k- and q- grids and , as shown in Figs. S5 S7. At last, we check whether the result is transferable to systems with larger unit cell where essentially smaller k-grids provide converged ground-state properties and q-grid, taken half of k-grid, sometime shrinks only to include the -point. To do that, we performed EPC calculations with 222 supercell of MgB2 which has 24 atoms per cell, and present the results in Fig. S8 and S9. Results confirm that q-grid as half as that of the k- grid can provide converged results even for larger systems. Instead of increasing coarse k- and q- grids, one can also utilize extremely large k- grid for interpolating EPC matrix element to achieve convergence, as shown in Figs. S7 and S10 respectively for MgB2 and AlB2. However, it also increases computational complexity for larger systems.
@@ XV Convergence ansatz
Our investigation revealed that a reasonable number of calculations did not converge, leading to inaccurate predictions, while utilizing a coarse k-grid from MP database and fine grid for interpolation only twice that of the k-grid [Table S7]. For example, AlB2 was predicted to be a superconductor (Tc 11 K) with the coarse grid [k-grid: 888, q-grid: 444, fine-k-mesh: 161616] from MP database, whereas a denser grid twice the size of the coarse grid corrected this inaccuracy and predicted AlB2 to be a nonsuperconductor, consistent with experimental observations [Fig. S7]. To identify such cases of convergence failure, we examined the variation of Tc with respect to and developed a simple ansatz based on the converged results of MgB2, as depicted in Fig. S13. The ansatz involves extracting Tc with a fixed broadening value (as in the case of MgB2) and estimating the decay parameter (A) in the exponential variation of Tc with as T. Unconverged results exhibit larger values of A, which decrease with a denser k-mesh.
@@
with also depends on . Despite of dependency of on EPC matrix () and phonon frequency () terms, the variation in is largely guided by the variation in or the Gaussian term G. Fig. S11(a) shows the variation of G with respect to smearing for different values of , assuming independent of . The effect of on G dictates the variation of and hence Tc, with depends on the materials as well as on size of the grid to compute double delta integration. A larger represents more states contributing towards the double delta summation within space around the Fermi-level spanned by . In other word, will increase with if the density of states has local maximum close to the Fermi-level, and the variation of EPC properties will be similar to the Gaussian plots corresponding to 0.05. For Niobium, the variation of EPC properties with respect to follows similar to blue-dashed curve corresponding to = 0.01 [3], whereas for MgB2, the variation follows the Gaussian corresponding to 0.01. For and , if the Gaussian function recovers the delta function, while for the Gaussian drops to zero similar to 0.01 cases. Our primary focus is more on exponentially decaying cases.
@@
where is the Debye cutoff energy and U is electron-phonon coupling potential. One can established the relation between or on as and with constant. Here parameter is a constant with a positive value if decreases with increasing and becomes negative otherwise. Here A and B are constants to be determined with A denoting the coefficient of exponential increase (for negative A) or decrease (for positive A) of with respect to . To determine the value of these constants, we perform linear-fit of vs for different k- grids for MgB2, which is frequently studied both theoretically as well as experimentally and often challenging to obtain the converged SC properties [7, 8, 9]. Furthermore, one can capture the behavior of Tc vs for a wide range of Tc using MgB2, as shown in Fig. S11(b). We utilize a smaller k-mesh of 664 and a slightly larger k′-grid of 886 for comparison [Fig. S11(b) and (c)].
@@
Fig. S11 (b) represents the variation of with respect to , while Fig. S11 (c) presents vs plot. It exhibits nearly linear behavior up to , i.e. = 0.37, after which it deviates from linearity. The linear-fit has optimal coefficient of determination score (R2-score) for on vs data for k(charge density and EPC)-q(Phonon and EPC)-2k(interpolating EPC matrix) grids (Fig. S11 (c)) [Table within the Fig. S11(c)]. Therefore, we chose in this work with R2-score of 0.9987. Besides different choice of leads to different values of A and B, it doesn’t have a significant role. For coarse grid, the Tc has a larger dependency on and shows larger exponential decay, compared to more converged calculations on denser grids. The exponential decay parameter “A” decreases from 19 to 10 with grids changing from coarse to dense one. This analysis suggests that A = 1213 can be used as a cutoff for this work, the calculations can be considered unconverged for A A, while the calculations can be considered converged for A A for accurate predictions. Fig. S11(d) shows the convergence of the Tc with respect to q-point mesh, keeping k-grid fixed at 121212. A q-point grid as half as that of the ground-state k-point grid is sufficient to provide the converged results for MgB2. Fig. S11 (e) represents the spectral functions for various grids with k=886. Denser grid (2k) slightly blue shifts the spectral function peaks at lower frequency range (15.5-16.5 THz), while peaks at higher frequency ranges (20-22.5 THz) remain unaffected. This results slight change in from 0.78 to 0.72 and Tc changing from 20 K to 15-16 K [Figs. S5-S7]. A of 0.78 agrees with previous theoretical value of 0.75 from anisotropic Migdal-Eliashberg calculation using 0.16 [5], 0.71 from fully anisotropic SCDFT [10], and 0.73 from previous isotropic Eliashberg approximation [11] at slightly larger 0.015 Ry. This indicates a k-grid obtained from MP database with Ry already provide converged results in the case of MgB2. However, it is not always the case for other materials. Based on these results, parameters = 0.01 Ry with seems to be reasonable choice for smearing with 8 8 6 to 16 16 12 k-grids for MgB2, and for other systems for the sake of comparison. These parameters could also depend on pseudopotentials. In order to achieve converged results, it is necessary to use denser k- and q-grids. Subsequently, a double-delta integration should be performed, selecting the results that correspond to the limit of . Other details convergence tests of MgB2 with respect to Brillouin-zone sampling are presented in Figs. S5-S10.
@@
In Fig. S12, we show two cases of YB12 and C2I2Y2 in which the coarse k- grids from Materials Project and fine k- grid only twice of that result to qualitatively inaccurate results, unusually high Tc for the former while predicting negligible Tc for the latter. With coarse grids, the exponential decay parameter of YB12 is around 38, much larger than the critical value estimated for MgB2 (A 12). Increasing k-mesh grids two times in each direction, Tc drops from 40.6 K to 1.67 K at 0.01 Ry and drops to almost zero for 0.01 Ry indicating non-superconductor, which is in good agreement with the experiment [12]. We found a couple of cases such as C2I2Y2 when first decreases and increases later with (we set A = 0 in Fig. S14). Such behavior hasn’t been observed in G(,) vs plots for a wide range of (Fig. 8(a) of the main text). The coarse grids result to a very low Tc of 0.97 K for C2I2Y2. However, the denser mesh corrects the behavior with an exponential decay parameter of 12.5 much closer to the reference value of MgB2. Also, the critical temperature Tc improves to 5.29 K agreeing much more strongly to the experiment [13]. Note that, this analysis only works for compounds with non-zero for a wide range of . For 1 K at Ry (computational details for choosing Ry), 0.1 K for Ry, and rapidly drops to zero for larger , we identify the compound to be non-superconducting with the value of A as zero (A = 0 A, already converged). To summarize the ansatz, we present a simple schematics of the process shown as in Fig. S13.
@@
We present a few qualitatively inaccurate predictions with A A and improved results with a denser grid in Table S7, while Table S8 shows unaffected results (qualitatively) with denser grids for the cases with A A. For example, AlB2 is predicted to be superconductors with a coarse grid, while it changes to a non-superconductor with a denser grid, consistent with the experimental prediction. Similarly, of superconductors such as CaNiBN, C2I2Y2, and C2Cl2Y2 change from below 1 K (non-superconductor) to a larger value, and show a better agreement with the experiment. On the contrary, the qualitative results remain unchanged for converged calculations represented by A A, regardless of their agreement with the experimental references.
XVI Application of convergence ansatz on known compounds of SuperCon